Consider a population consisting of 1,2,3,4 and 5. Supposed samples of size 2 are drawn from this population. describe the sampling distribution of the sample means.

  • What is the mean and variance of the sampling distribution of the sample means?
  • Compare these values to the mean and variance of the population.
  • Draw the histogram of the sampling distribution of the population mean.
1
Expert's answer
2021-03-26T17:40:29-0400

Here population size : N = 5

and which are 1,2,3,4 and 5

And we have to draw a sample of size 2

So, there are "^{N}C_n" possible samples

that is, "^{5}C_2= 10"

Mean of population "(\mu)" = "\dfrac{1+2+3+4+5}{5}=3"


Variance of population "(\sigma^2)=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{4+1+0+1+4}{5}=2"

So all the possible samples are:

(1,2),(1,3),(1,4),(1,5),(2,3),(2,4),(2,5),(3,4),(3,5) and (4,5)


Mean of sampling distribution "(\mu_{\bar{x}})=\mu=3"


The variance of sampling distribution "(\sigma^2_{\bar{x}})= \dfrac{\sigma^2}{N}=\dfrac{4}{5}=0.8"


So mean of population = mean of sample = 3

Variance of population = 2 and variance of sample = 0.8




Histogram of the mean of population:




Histogram of the mean of sampling: